# Modeling Ultra-High-Energy Cosmic Rays propagation using the input from Configuration Interaction Shell Model

### v1

@@ IV Application to UHECR propagation
To describe the changes in abundance of the heavy nuclei as a result of the interaction of the UHECR with the CMB, a nuclear reaction network including all interactions of interest must be considered. The chosen set of nuclear species are coupled by a system of differential equations corresponding to all the reactions affecting each nucleus, i.e. mainly photodisintegrations and $\beta$-decays. The rate of change of the number density $N_{Z,A}$ of a nucleus $(Z,A)$ with charge number $Z$ and mass number AA can be written as
@@
where $\lambda_{\beta}^{Z,A}$ is the $\beta$-decay rate of nucleus $(Z,A)$ and $\lambda_{\gamma,x}^{Z,A}$ its photodisintegration rate followed by the emission of a single neutron ($x=n$), proton ($x=p$) or $\alpha$-particle ($x=\alpha$) or the emission of multiple particles such as $2n$, $2p$, $2\alpha$, npnp, $\dots$, including all open channels for a given photon energy distribution.
@@ V Conclusion
It should be noted that nuclei up to 56Fe are expected to be present in the UHECR source; thus the CI-SM predictions of this work covered only $\sim 50\%$ of nuclei necessary for the complete modeling of UHECR. Based on the previous CI-SM calculations of [59], we are currently developing a suitable approach to also address dipole response of pfpf-shell nuclei. The current diagonalization limit of $\sim 10^{10}$ will allow us to compute the $E1$ strength function of pfpf-shell nuclei up to 50Fe, without any further approximation, and thus to fill in further the existing gap of CI-SM predictions of relevance. While here we examined how the PSF model may affect the UHECR interaction with the CMB, it was demonstrated [60] that the impact of nuclear model uncertainties is potentially larger in environments with non-thermal radiation fields, e.g. in the possible astrophysical sources of UHECR such as the gamma-ray bursts. Such sensitivity studies could be performed in the future with TALYS cross sections based on the microscopic CI-SM predictions.

### v2

@@ IV Application to UHECR propagation
To describe the changes in abundance of the heavy nuclei as a result of the interaction of the UHECR with the CMB, a nuclear reaction network including all interactions of interest must be considered. The chosen set of nuclear species are coupled by a system of differential equations corresponding to all the reactions affecting each nucleus, i.e. mainly photodisintegrations and $\beta$-decays. The rate of change of the number density $N_{Z,A}$ of a nucleus $(Z,A)$ with charge number $Z$ and mass number $A$ can be written as
@@
where $\lambda_{\beta}^{Z,A}$ is the $\beta$-decay rate of nucleus $(Z,A)$ and $\lambda_{\gamma,x}^{Z,A}$ its photodisintegration rate followed by the emission of a single neutron ($x=n$), proton ($x=p$) or $\alpha$-particle ($x=\alpha$) or the emission of multiple particles such as $2n$, $2p$, $2\alpha$, $np$, $\dots$, including all open channels for a given photon energy distribution.
@@ V Conclusion
It should be noted that nuclei up to 56Fe are expected to be present in the UHECR source; thus the CI-SM predictions of this work covered only $\sim 50\%$ of nuclei necessary for the complete modeling of UHECR. Based on the previous CI-SM calculations of [59], we are currently developing a suitable approach to also address dipole response of $pf$-shell nuclei. The current diagonalization limit of $\sim 10^{10}$ will allow us to compute the $E1$ strength function of $pf$-shell nuclei up to 50Fe, without any further approximation, and thus to fill in further the existing gap of CI-SM predictions of relevance. While here we examined how the PSF model may affect the UHECR interaction with the CMB, it was demonstrated [60] that the impact of nuclear model uncertainties is potentially larger in environments with non-thermal radiation fields, e.g. in the possible astrophysical sources of UHECR such as the gamma-ray bursts. Such sensitivity studies could be performed in the future with TALYS cross sections based on the microscopic CI-SM predictions.